OpenStudy (anonymous):

Find the solution of the differential equation dP/dt=2(sqrt(Pt)), that satisfies the given initial condition P(1)=2

OpenStudy (anonymous):

Is that \[\frac{dP}{dt}=2\sqrt{Pt}~~?\]

OpenStudy (anonymous):

If it is, then the equation is separable: \[\frac{dP}{dt}=2\sqrt{Pt}~~\iff~~\frac{dP}{\sqrt P}=2\sqrt t~dt\]

OpenStudy (anonymous):

that is correct. Where do i go from there? Thanks for the help!

OpenStudy (anonymous):

Recall that \(\sqrt x=x^{1/2}\) and \(\dfrac{1}{\sqrt x}=\dfrac{1}{x^{1/2}}=x^{-1/2}\). It's power rule from there.

OpenStudy (anonymous):

\[P^{-1/2}~dP=2t^{1/2}~dt\] Integrating both sides yields \[\frac{P^{1/2}}{1/2}=2\frac{t^{3/2}}{3/2}+C~~\iff~~2P^{1/2}=\frac{4}{3}t^{3/2}+C\]

OpenStudy (anonymous):

Solve for \(C\) with the given condition, \(P(1)=2\): \[2(2^{1/2})=\frac{4}{3}(1^{3/2})+C~~\Rightarrow~~2^{3/2}=\frac{4}{3}+C~~\Rightarrow~~C=\cdots\] then plug it into the general solution to get the particular one.

OpenStudy (anonymous):

so is the answer [P=\frac{ 1 }{ 2 }(\frac{ 4 }{ 3 }t ^{3/2}+\sqrt{3}/2)^{2}\]

OpenStudy (anonymous):

\[\begin{align*}2P^{1/2}&=\frac{4}{3}t^{3/2}+\underbrace{2^{3/2}-\frac{4}{3}}_{C}\\\\ P^{1/2}&=\frac{2}{3}t^{3/2}+2^{1/2}-\frac{2}{3}\\\\ P&=\left(\frac{2}{3}t^{3/2}+\sqrt2-\frac{2}{3}\right)^2 \end{align*}\] I'd stop there, but you could expand if you like